OpenStudy (anonymous):

Convert these unlike fractions to equivalent like fractions and add them. You must use the LCD to get the answer correct. If possible, reduce the final sum 1/4+5/12= 1/4= 5/12=

OpenStudy (anonymous):

1/4=5/12= ? and ?